If 73.5 mol of an ideal gas occupies 72.5 L at 73.00 °C, what is the pressure of the gas? P-
Answer #1

For the given ideal gas:

Volume( V ) = 72.5L

Temperature( T ) = 73°C = (273 + 73) K = 346K

Moles (n) = 73.5 mol

Pressure ( P ) = ?

We know that for an ideal gas -

PV = nRT where R = 0.0821 L-atm /K /mol is the universal gas gas constant.

=> P = nRT/V = ( 73.5 × 0.0821 × 346 ) / 72.5 = 28.798 atm